In ∆abc, s-a/∆=1/8, s-b/∆=1/12, s-c/∆=1/24 , find the value of b?
 

Grade:11

1 Answers

Spacspade
6 years ago
I’m assuming your question is (s-a)/(delta)=1/8 and so on..
So multiply the 3 equations and you will get (delta).s=8*12*24
then find whatever you require!